OpenStudy (daisiablis11):

I'm a little confused on how (1/x)/(1/2sqrt(x)) = 2/sqrt(x)

satellite73 (satellite73):

dividing by a fraction is the same as multiplying by the reciprocal

satellite73 (satellite73):

i.e. \[\frac{\frac{1}{x}}{\frac{1}{2\sqrt2}}\] \[=\frac{1}{x}\times \frac{2\sqrt2}{1}\]

satellite73 (satellite73):

dang typo \[=\frac{1}{x}\times \frac{2\sqrt{x}}{1}\]

OpenStudy (mhchen):

\[\frac{\frac{1}{x}}{\frac{1}{2\sqrt{x}}} = \frac{1}{x} * \frac{2\sqrt{x}}{1} = \frac{2\sqrt{x}}{x} *(\frac{\sqrt{x}}{\sqrt{x}})=\frac{2x}{x \sqrt{x}} = \frac{2}{\sqrt{x}} \]

OpenStudy (daisiablis11):

Thank you both.